If each of the data set is increased by 10% what happens to the value of the
Mean
Mean deviation
Standard deviation
1
Expert's answer
2020-04-17T17:25:30-0400

Let data points as "x_i" and number of data points as n.

After increase every data"(x_i)" point by 10%,new data points"(y_i)" are,

"y_i=x_i+x_i*0.1=1.1*x_i"


mean of data set before increment"(\bar{x})" ,

"\bar{x}=\frac{\sum\limits_{i=1}^{n}{x_i}}{n}\\"

mean of the data set after increment"(\bar{y})" ,

"\bar{y}=\frac{\sum\limits_{i=1}^{n}{y_i}}{n}\\
\bar{y}=\frac{\sum\limits_{i=1}^{n}{1.1*x_i}}{n}\\
\bar{y}=1.1\frac{\sum\limits_{i=1}^{n}{x_i}}{n}\\
\bar{y}=1.1\bar{x}"

Therefore mean is also increment by 10% when data set increment by 10%.


mean deviation of the data set before increment"(M_x),"

"M_x=\frac{\sum \limits_{i=1}^{n}{\text{\textbar} x_i-\bar{x}\text{\textbar}}}{n}"

mean deviation of the data set before increment"(M_y),"

"M_y=\frac{\sum \limits_{i=1}^{n}{\text{\textbar} y_i-\bar{y}\text{\textbar}}}{n}\\
M_y=\frac{\sum \limits_{i=1}^{n}{\text{\textbar}1.1* x_i-1.1*\bar{x}\text{\textbar}}}{n}\\
M_y=\frac{1.1*\sum \limits_{i=1}^{n}{\text{\textbar} x_i-\bar{x}\text{\textbar}}}{n}\\
M_y=1.1*M_x"

Therefore mean deviation is also increment by 10% when data set increment by 10%.


variance of the data set before increment"(\sigma^2_x)"

"\sigma^2_x=\frac{\sum\limits_{i=1}^{n}{(x_i-\bar{x})^2}}{n}=\frac{\sum\limits_{i=1}^{n}{x_i^2}}{n}-(\bar{x})^2\\"

variance of the data set after increment"(\sigma^2_y)"

"\sigma^2_y=\frac{\sum\limits_{i=1}^{n}{y_i^2}}{n}-(\bar{y})^2\\
\sigma^2_y=\frac{\sum\limits_{i=1}^{n}{(1.1*x_i)^2}}{n}-(1.1*\bar{x})^2\\
\sigma^2_y=(1.1)^2*\frac{\sum\limits_{i=1}^{n}{(x_i)^2}}{n}-(1.1)^2*(\bar{x})^2\\
\sigma^2_y=(1.1)^2*\sigma^2_x\\
\sigma^2_y=1.21*\sigma^2_x"

standard deviation,

"\sigma_y=1.1*\sigma_x"

Therefore standard deviation is also increment by 10% when data set increment by 10%.